OpenStudy (anonymous):

If f(x) = 5x – 3, then what is f–1(x) ? I really don't get this :/

OpenStudy (anonymous):

the -1 is a ^-1 sorry

OpenStudy (anonymous):

f^-1

OpenStudy (anonymous):

f^-1(x) is the inverse function of f(x) y = 5x - 3 x = 5y - 3 x + 3 = 5y y = (x+3)/5 f^-1(x) = (x+3)/5

OpenStudy (anonymous):

Thank you :)